OpenStudy (anonymous):

log((3x^2)/12) I needed to know if this is the answer when you expand (log 3 - 2logx - log 12)

OpenStudy (anonymous):

It is not the answer

OpenStudy (anonymous):

Log(3) + 2log(x) - log(12)

OpenStudy (anonymous):

remember that log(x) + log(b) = log(xb) and log(x) - Log(b) = log(xb) also -1Log(x) = Log(1/x) because x^(-1) = 1/x

OpenStudy (anonymous):

remember when you see just a negative symbol there is a -1 there